OpenStudy (anonymous):

Find the missing value. 3:? :: 10:50

OpenStudy (anonymous):

let the number be x then 3:x:: 10:50 3/x=10/50 x=(3x50)/10=15

OpenStudy (anonymous):

Thank you for ur help and how to show the work :)